(3x2 +y2)dy = (-x2- 3y2)dx


1
Expert's answer
2021-02-16T12:35:28-0500

Solution.

"(3x^2+y^2)dy=(-x^2-3y^2)dx."

This is a homogeneous nonlinear differential equation of the first order. Divide both parts of the equation by "dx:"

"(3x^2+y^2)\frac{dy}{dx}=-x^2-3y^2, \text{or } (3x^2+y^2)y'=-x^2-3y^2."

Make a replacement

"u(x)=\frac{y(x)}{x}," from here "y(x)=xu(x)," and "y'=u(x)+xu'(x)."

Substitute the obtained expressions in our equation:

"(3x^2+x^2u^2)(u+xu')+x^2+3x^2u^2=0,"

"x^3u^2u'+3x^3u'+x^2u^3+3x^2u^2+3x^2u+x^2=0,"

"x^3u'(u^2+3)+x^2(u^3+3u^2+3u+1)=0."

Divide both parts of the equation by "u^3+3u^2+3u+1" and write the equation in the form:


"\frac{(u^2+3)u'}{u^3+3u^2+3u+1}=-\frac{1}{x}."

Thus we obtain a differential equation with dividing terms. Let's solve it:


"\frac{dx(u^2+3)\frac{du}{dx}}{u^3+3u^2+3u+1}=-\frac{dx}{x},"

"\frac{du(u^2+3)}{u^3+3u^2+3u+1}=-\frac{dx}{x},"

"\int\frac{u^2+3}{u^3+3u^2+3u+1}du=\int-\frac{1}{x}dx,"

"\frac{2u}{u^2+2u+1}+\ln(u+1)=-\ln{x}+C,"

where C - some constant.

Returning to the replacement we obtain the solution of the given equation:


"\frac{2y(x)}{x(1+\frac{2y(x)}{x}+\frac{y^2(x)}{x^2})}+\ln(1+\frac{y(x)}{x})=-\ln{x}+C,"

or


"\frac{2xy(x)}{(y(x)+x)^2}+\ln(1+\frac{y(x)}{x})=-\ln{x}+C."

Answer.

"\frac{2xy(x)}{(y(x)+x)^2}+\ln(1+\frac{y(x)}{x})=-\ln{x}+C."